A beaker contains 11.30 M osimium (lll) fluoride in 0.0673 liters of solution. How many grams of osimium (lll) fluoride are dissolved in this solution?


1
Expert's answer
2021-05-26T07:12:37-0400

It is given that:


Molarity (M) = 11.30 M

Volume (V) = 0.0673 liters

Now :


"Molarity(M)=\dfrac{moles}{Volume(liters)}"


"11.30=\dfrac{moles}{0.0673}"


So Moles of osimium(III) fluoride "=11.30\times0.0673=0.76moles"

Now mass of osimium(III) fluoride "=moles\times molar mass"

"=0.76\times304.2=231.192 gm"